OpenStudy (anonymous):

what is the most appropriate method for the equation 4x^2 - 9 = 0 ? the method of factoring or the Extracting the square roots?

OpenStudy (solomonzelman):

\(\Large\color{darkgreen}{ \rm 4x^2 - 9 = 0 }\) \(\Large\color{darkgreen}{ \rm (2x)^2 - 3^2 = 0 }\) \(\Large\color{darkgreen}{ \rm (2x-3)(2x+3) = 0 }\) finish it....

OpenStudy (anonymous):

what is the best method of it ? by factoring ? or the other one?

OpenStudy (triciaal):

extracting the root difference of 2 squares 4x^2 =9 2x = +/- 3 x = +/- 3/2

OpenStudy (solomonzelman):

well either way. Factoring a²-b² into (a-b)(a+b) , or adding 9 to both sides and completing the square. Same thing :)

OpenStudy (triciaal):

@SolomonZelman don't get me wrong It's just the way the question is worded. Which is more appropriate for this particular situation. for all purposes could also use the quadratic formula.

OpenStudy (solomonzelman):

well, if you mean that I would get offended.. not at all. I do like your method better though. I myself a fan of completing the square.

OpenStudy (triciaal):

Thanks. my preference is looking at the numbers and if I cannot easily identify factors then use the formula. fastest for me. First things first, what is the question asking.